Mrs. Grasso’s eleventh-grade class has 12 boys and 14 girls. In how many ways can
at least 2 boys be selected for a committee of 3 students?
(A) 66 (B) 286 (C) 924 (D) 1,144 (E) 4,004

Answer by sudhanshu_kmr(1152)


In each committee should have 3 student.

no. of ways to select 2 boys and 1 girl = 12C2 * 14C1 = 66*14 =

no. of ways to select 3 boys = 12C3 = 220


total no. of ways to select at least 2 boys for committee of 3 = 924 + 220 = 1144

answer is 1144